# LLM Unlearning with LLM Beliefs

### v1

@@ E.4 Hyperparameter Settings of Unlearning Methods
In Appendix F.5.1, we present a sample of hyperparameter selection on TOFU 10% with LLaMA-3.1-8B.
@@ Appendix F Additional Results
This appendix section provides complementary results to support the main text. Specifically, Appendix F.1 offers further qualitative examples of GA and NPO to highlight their limitations; Appendix F.2 details the prompt design used in our proposed LaaJ evaluation; Appendix F.3 presents additional results on the MUSE-Books benchmark; Appendix F.4 provides illustrative outputs of different unlearning methods; Appendix F.5 reports extended analyses including hyperparameter sensitivity and the effect of alternative unlearning losses in BS-S; and Appendix F.6 summarizes training time comparisons across methods. Together, these results offer a more comprehensive view of both prior baselines and our proposed framework.
@@ F.1 Additional Examples on Failure Cases
We here provide additional qualitative examples under the 10% TOFU forget setting with LLaMA-3.1-8B-Instruct to further illustrate typical failure modes of GA- and NPO-based unlearning. These cases show that GA fails by collapsing the entire model, while NPO fails by allowing semantic rephrasings to persist.
@@ F.2 Prompt Design of Proposed LaaJ Evaluation
Here, we provide the detailed prompt design for our LaaJ evaluation introduced in Section 3.1: Similarity and Naturalness, which assess whether unlearning is both semantically thorough and linguistically fluent. In practice, we adopt Gemini 2.5 Flash (Comanici et al., 2025) for evaluation, while other advanced models such as ChatGPT 5 or Claude 4.1 can also be employed.
@@
F.3 Results on MUSE-Books
Table 5 reports results on MUSE-Books with LLaMA-2-7B-chat. While SimNPO achieves the highest UtilPres, its forgetting is far from complete, with both VerbMem and KnowMem remaining non-negligible. In contrast, our bootstrapping methods drive both forgetting scores down to zero, ensuring thorough removal of targeted knowledge. Among these methods, BS-S attains the strongest utility preservation (UtilPres 0.3854), slightly outperforming BS-T (0.3842) and clearly surpassing other approaches that also reduce memorization to zero. This highlights that our framework not only achieves complete forgetting but also retains more utility compared with competing baselines.
Table 5: Performance with retain regularization on MUSE-Books with LLaMA-2-7B-chat.
@@ F.4 Qualitative Examples of Unlearning Results
In this subsection, we present qualitative examples of model responses after applying different unlearning methods on TOFU under the 10% forget setting with LLaMA-3.2-1B-Instruct. These examples allow us to directly compare how various approaches behave once unlearning is applied. For each case, we highlight the limitations of baseline methods such as GradDiff, NPO, and RMU, and contrast them with our bootstrapping approaches (BS-T and BS-S).
Example 1 (Hsiao Yun-Hwa’s writing language). Baseline methods such as GradDiff and NPO still explicitly mention “English,” thereby leaking the exact sensitive knowledge that should have been removed. RMU and SimNPO generate partially distorted responses, but they either retain fragments of the ground-truth or introduce implausible additions. In contrast, BS-T reframes the answer by providing an alternative language (Taiwanese dialect), effectively breaking the direct link to the ground-truth, while BS-S generates a longer, more diverse narrative that, despite some noise, clearly avoids reproducing the original sensitive information.
@@ F.5.1 Hyperparameter Analysis
We conduct ablation studies on the hyperparameters of BS-T and BS-S using the aggregate score (Agg.) as the criterion for parameter selection. Figure 5 reports results under the 10% forget setting on TOFU with LLaMA-3.1-8B. For BS-T, the two hyperparameters are the interpolation weight $\lambda_{\rm BST}$ and the neighborhood size $k$. For BS-S, the hyperparameters are the number of sampled sequences $N$ and the interpolation weight $\lambda_{\rm BSS}$.
From Figure 5(a), we observe that BS-T is particularly sensitive to $\lambda_{\rm BST}$. The aggregate score peaks at $\lambda=0.2$ and $k=10$, reaching $0.63$. Increasing or decreasing $\lambda_{\rm BST}$ beyond this point results in a significant drop in performance, indicating that an overly small weight reduces the suppression of high-likelihood alternatives, while an overly large weight excessively penalizes the model and harms utility. In contrast, the effect of $k$ is relatively modest: changing the neighborhood size from $5$ to $50$ only slightly alters the performance, confirming that BS-T mainly relies on the balance set by $\lambda_{\rm BST}$. Notably, setting $\lambda_{\rm BST}=0$ recovers GradDiff, providing a useful reference baseline.
Turning to Figure 5(b), we find that BS-S behaves differently. The parameter $N$ improves performance with diminishing returns: raising $N$ from $1$ to $3$ yields notable gains, but further increases provide only marginal improvements. On the other hand, $\lambda_{\rm BSS}$ again plays a crucial role, producing large variations across different values. Notably, the best performance is achieved at $N=10$ and $\lambda_{\rm BSS}=0.6$, with an aggregate score of $0.65$. However, we remark that setting $\lambda_{\rm BSS}=0$ effectively reduces BS-S to BS-T, highlighting the role of sequence-level bootstrapping. Moreover, the memory footprint of BS-S grows linearly with $N$, and beyond $N=5$ a single 8080G GPU cannot fit the training, leading to out-of-memory errors. Since the empirical gains also saturate, we restrict our hyperparameter search up to $N=5$ in practice.
@@
Figure 5: Aggregate score (Agg.) of BS-T and BS-S under different combinations of hyperparameters on TOFU 10% forget setting with LLaMA-3.1-8B.
@@ F.5.2 Unlearning Loss in BS-S
Table 6: Ablation on BS-S by replacing its unlearning loss on TOFU 10% with LLaMA-3.1-8B.
@@
w/ BST (BS-S) $0.64$ $0.58$ $0.71$
We further investigate the effect of different unlearning losses when plugged into the BS-S framework. As shown in Table 6, replacing the loss with GA, NPO, or WGA still benefits from the sequence-level bootstrapping design, consistently improving the aggregate score compared to their standalone counterparts in Table 1. Among them, BS-S with BS-T loss achieves the best overall performance, attaining the highest Agg. and Mem. while preserving utility competitively. This confirms that the proposed BS-S formulation not only generalizes across different unlearning objectives, but also works most effectively with the BS-T loss.
@@ F.6 Time Consumption
Figure 6 reports the average per-epoch training time on the TOFU 10% forget setting with LLaMA-3.2-1B-Instruct, using a single NVIDIA A6000 GPU and batch size of 32. Our bootstrapping methods are efficient compared with heavy approaches such as GRU (Wang et al., 2025d). BS-T is relatively fast since it only modifies the loss function without changing the data pipeline, leading to a runtime close to lightweight baselines like RMU and WGA. By contrast, BS-S is slower because it requires additional model sampling and data augmentation at each step, which increases overhead. Nevertheless, the extra cost of BS-S remains acceptable given that it achieves the best unlearning–utility trade-off. These results show that our methods combine strong effectiveness with manageable training cost under practical settings.
Figure 6: Training time (s) per epoch on TOFU 10% forget setting with LLaMA-3.2-1B.

### v2

@@ E.4 Hyperparameter Settings of Unlearning Methods
In Appx. F.5.1, we present a sample of hyperparameter selection on TOFU 10% with Llama 3.1 8B.
@@ Appendix F Additional Results
This appendix section provides complementary results to support the main text. Specifically, §F.1 offers further qualitative examples of GA and NPO to highlight their limitations; §F.2 details the prompt design used in our proposed LaaJ evaluation; §F.3 presents additional results on the MUSE-Books benchmark; §F.4 provides illustrative outputs of different unlearning methods; §F.5 reports extended analyses including hyperparameter sensitivity and the effect of alternative unlearning losses in BS-S; and §F.6 summarizes training time comparisons across methods. Together, these results offer a more comprehensive view of both prior baselines and our proposed framework.
@@ F.1 Additional Examples on Failure Cases
We here provide additional qualitative examples under the 10% TOFU forget setting with Llama 3.1 8B-Instruct to further illustrate typical failure modes of GA- and NPO-based unlearning. These cases show that GA fails by collapsing the entire model, while NPO fails by allowing semantic rephrasings to persist.
@@ F.2 Prompt Design of Proposed LaaJ Evaluation
Here, we provide the detailed prompt design for our LaaJ evaluation introduced in §3.1: Similarity and Naturalness, which assess whether unlearning is both semantically thorough and linguistically fluent (Zhou et al., 2024; 2025; 2026). In practice, we adopt Gemini 2.5 Flash (Comanici et al., 2025) for evaluation, while other advanced models such as GPT-5.2 or Claude 4.6 can also be employed.
@@
F.3 Results on MUSE
Table 4: Performance with retain regularization on MUSE-News with Llama 2 7B-Chat.
Forget Retain
Method VerbMem KnowMem UtilPres
Original $0.5789$ $0.6443$ $0.5552$
Retrain $0.2016$ $0.3170$ $0.5602$
GradDiff $0.3249$ $0.3481$ $0.4603$
NPO $0.2914$ $0.3290$ $0.4651$
RMU $0.3861$ $0.5088$ $0.4962$
SimNPO $0.4608$ $0.6043$ $0.5010$
WGA $0.3713$ $0.5732$ $0.4782$
BS-T (Ours) $0.2837$ $0.3278$ $0.4602$
BS-S (Ours) $0.2713$ $0.3250$ $0.4774$
Tab. 4 reports results on MUSE-News. Our methods achieve the lowest forget scores, with BS-S reducing VerbMem/KnowMem to 0.2713/0.3250 compared to 0.2914/0.3290 for NPO and 0.3861/0.5088 for RMU. At the same time, BS-S attains higher UtilPres (0.4774) than most baselines, showing a better balance between forgetting and retention. BS-T also improves forgetting effectiveness (0.2837/0.3278) while maintaining competitive utility (0.4602). These results highlight that explicitly unlearning both original harmful targets and model beliefs enables our methods to deliver more thorough forgetting with minimal loss in retained performance.
Table 5: Performance with retain regularization on MUSE-Books with Llama 2 7B-Chat.
@@ F.3 Results on MUSE
Tab. 5 reports results on MUSE-Books. While SimNPO achieves the highest UtilPres, its forgetting is far from complete, with both VerbMem and KnowMem remaining non-negligible. In contrast, our bootstrapping methods drive both forgetting scores down to zero, ensuring thorough removal of targeted knowledge. Among these methods, BS-S attains the strongest utility preservation (UtilPres 0.3854), slightly outperforming BS-T (0.3842) and clearly surpassing other approaches that also reduce memorization to zero. This highlights that our framework not only achieves complete forgetting but also retains more utility compared with competing baselines.
@@ F.4 Qualitative Examples of Unlearning Results
In this subsection, we present qualitative examples of model responses after applying different unlearning methods on TOFU under the 10% forget setting with Llama 3.2 1B-Instruct. These examples allow us to directly compare how various approaches behave once unlearning is applied. For each case, we highlight the limitations of baseline methods such as GradDiff, NPO, and RMU, and contrast them with our bootstrapping approaches (BS-T and BS-S).
Example 1 (Hsiao Yun-Hwa’s writing language). Baseline methods such as GradDiff and NPO still explicitly mention “English”, thereby leaking the exact sensitive knowledge that should have been removed. RMU and SimNPO generate partially distorted responses, but they either retain fragments of the ground-truth or introduce implausible additions. In contrast, BS-T reframes the answer by providing an alternative language (Taiwanese dialect), effectively breaking the direct link to the ground-truth, while BS-S generates a longer, more diverse narrative that, despite some noise, clearly avoids reproducing the original sensitive information.
@@ F.5.1 Hyperparameter Analysis
We conduct ablation studies on the hyperparameters of BS-T and BS-S using the aggregate score (Agg.) as the criterion for parameter selection. Fig. 5 reports results under the 10% forget setting on TOFU with Llama 3.1 8B. For BS-T, the two hyperparameters are the interpolation weight $\lambda_{\rm BST}$ and the neighborhood size $k$. For BS-S, the hyperparameters are the number of sampled sequences $N$ and the interpolation weight $\lambda_{\rm BSS}$.
From Fig. 5(a), we observe that BS-T is particularly sensitive to $\lambda_{\rm BST}$. The aggregate score peaks at $\lambda=0.2$ and $k=10$, reaching $0.63$. Increasing or decreasing $\lambda_{\rm BST}$ beyond this point results in a significant drop in performance, indicating that an overly small weight reduces the suppression of high-likelihood alternatives, while an overly large weight excessively penalizes the model and harms utility. In contrast, the effect of $k$ is relatively modest: changing the neighborhood size from $5$ to $50$ only slightly alters the performance, confirming that BS-T mainly relies on the balance set by $\lambda_{\rm BST}$. Notably, setting $\lambda_{\rm BST}=0$ recovers GradDiff, providing a useful reference baseline.
Turning to Fig. 5(b), we find that BS-S behaves differently. The parameter $N$ improves performance with diminishing returns: raising $N$ from $1$ to $3$ yields notable gains, but further increases provide only marginal improvements. On the other hand, $\lambda_{\rm BSS}$ again plays a crucial role, producing large variations across different values. Notably, the best performance is achieved at $N=10$ and $\lambda_{\rm BSS}=0.6$, with an aggregate score of $0.65$. However, we remark that setting $\lambda_{\rm BSS}=0$ effectively reduces BS-S to BS-T, highlighting the role of sequence-level bootstrapping. Moreover, the memory footprint of BS-S grows linearly with $N$, and beyond $N=5$ a single 80G GPU cannot fit the training, leading to out-of-memory errors. Since the empirical gains also saturate, we restrict our hyperparameter search up to $N=5$ in practice.
@@
Figure 5: Aggregate score (Agg.) of BS-T and BS-S under different combinations of hyperparameters on TOFU 10% forget setting with Llama 3.1 8B.
@@ F.5.2 Unlearning Loss in BS-S
Table 6: Ablation on BS-S by replacing its unlearning loss on TOFU 10% with Llama 3.1 8B.
@@
w/ BS-T (BS-S) $0.64$ $0.58$ $0.71$
We further investigate the effect of different unlearning losses when plugged into the BS-S framework. As shown in Tab. 6, replacing the loss with GA, NPO, or WGA still benefits from the sequence-level bootstrapping design, consistently improving the aggregate score compared to their standalone counterparts in Tab. 1. Among them, BS-S with BS-T loss achieves the best overall performance, attaining the highest Agg. and Mem. while preserving utility competitively. This confirms that the proposed BS-S formulation not only generalizes across different unlearning objectives, but also works most effectively with the BS-T loss.
@@ F.6 Time Consumption
Fig. 6 reports the average per-epoch training time on the TOFU 10% forget setting with Llama 3.2 1B-Instruct, using a single NVIDIA A6000 GPU and batch size of 32. Our bootstrapping methods are efficient compared with heavy approaches such as GRU (Wang et al., 2025d). BS-T is relatively fast since it only modifies the loss function without changing the data pipeline, leading to a runtime close to lightweight baselines like RMU and WGA. By contrast, BS-S is slower because it requires additional model sampling and data augmentation at each step, which increases overhead. Nevertheless, the extra cost of BS-S remains acceptable given that it achieves the best unlearning–utility trade-off. These results show that our methods combine strong effectiveness with manageable training cost under practical settings.
Figure 6: Training time (s) per epoch on TOFU 10% forget setting with Llama 3.2 1B.